### v1

@@ Abstract
In this work, we present results for discrimination of neutron and $\gamma$ events using a plastic scintillator detector with pulse shape discrimination capabilities. Machine learning (ML) algorithms are used to improve the discriminatory power between neutron and $\gamma$ events at lower energy ranges which otherwise are not addressed by the conventional pulse shape discrimination techniques. The use of a multilayer perceptron with Bayesian inference (MLPBNN) and support vector machine (SVM) algorithms are studied using the recorded waveforms from the detector. Three variables are constructed for the input of ML algorithms, which captures the essence of the differences in the head and tail part of the neutron and $\gamma$ waveforms. The training and the testing of the ML algorithms are done using an AmBe neutron source. In the lower energy region, the results obtained from the ML predictions are compared with the results obtained from a time of flight (ToF) technique to benchmark the overall performance of the ML algorithms. A reasonable agreement is observed between the results obtained from ML algorithm and the ToF experiment in the studied energy range. The MLPBNN gives better discriminatory power for the neutron and $\gamma$ events than the SVM algorithm.
@@ 1 Introduction
Pulse Shape Discrimination (PSD) is a widely used technique for the particle identification in nuclear and particle physics experiments [1, 2, 3], as well as in applications such as reactor environment monitoring [4] and the national security [5]. Among the various PSD methodologies, the charge comparison method (CCM) is one of the most reliable techniques for categorizing different types of ionizing radiation.
@@
Liquid organic scintillators are well-suited for PSD applications due to their well-defined scintillation time profiles for different particle interactions. The CCM effectively enhances these differences, making it a go-to method for particle identification in particle physics experiments. In particular, the use of this technique in reactor based neutrino detection experiments and rare event decay searches [6, 7, 8], where precise discrimination of neutron and $\gamma$ signals are required, to enhance the signal to background strength. In a very short baseline reactor based experiment, where both natural radioactivity and neutron-induced backgrounds (including fast and thermal neutrons) are prevalent, CCM provides a means of separating neutron-induced events from the actual signal of interest. Recently, developments in the plastic scintillator domain are being carried out to improvise the PSD properties [9, 10]. Easy to handle, large scale production and modularity are some of the advantages of plastic scintillators and can be efficiently used as core detector for the anti-neutrino detection [11, 12] and veto systems for muons and neutrons.
@@
Despite its advantages, the performance of CCM degrades significantly at lower energies, posing a major challenge for experiments demanding high background rejection sensitivities. The fast neutrons, in a typical reactor based very short baseline experiments (few meter from the reactor core), can mimic the prompt signal from positron events. This process leads to a potential misidentification of genuine neutrino events, even with the optimal shielding configurations [13, 14]. Therefore, a comprehensive understanding of both fast and thermal neutron backgrounds is vital for accurate event classification to improve the fidelity of neutrino and rare event search experiments. Apart from the particle physics experiments PSD technique is also extensively used in low energy nuclear physics experiments, especially in prompt fission neutron multiplicity measurements [15], can also benefit from these PSD based plastic scinitillators for neutron and $\gamma$ events.
In this study, supervised Machine Learning (ML) algorithms are explored to address the inefficiency of PSD CCM in the lower energy regime. Two approaches are investigated, a Multi-Layer Perceptron Bayesian Neural Network (MLPBNN) and a Support Vector Machine (SVM). These algorithms are trained and evaluated using data from a plastic scintillator detector, with PSD capabilities, to assess their ability to distinguish between particle types more effectively at lower energies. To check the effectiveness of the considered ML algorithms, a time-of-flight experiment is also performed, where low energy neutron and $\gamma$ events are distinguished and compared with the results obtained from ML algorithms.
@@ 2 Experimental Setup
A Plastic Scintillator (PS), EJ276D with PSD property from Eljen technology [16], is used to discriminate the signals from neutron and $\gamma$ events. The scintillator is of dimension $\phi$51$\times$51 mm and has a maximum emission wavelength of 425 nm. The detector is optically coupled with a $\phi$51 mm diameter Hamamatsu R7724, 10 stage photo-multiplier tube (PMT) of borosilicate glass window, which is operated at the negative bias voltage of 1200 V. The PSD ability of this particular PS are studied before and a decent neutron and $\gamma$ discrimination is obtained for higher energies [17, 18]. The data acquisition system (DAQ), used to measure the anode signal from the PS, consists of a 8-channel CAEN V1751 digitizer [19]. On-board digitization of the anode signal from the PMT is done using a leading edge discrimination technique. Triggered data is read by a Flash ADC of 10-bit resolution at 1 GSPS sampling rate. Raw waveforms are recorded and transferred to the computer via optical link for the detailed offline data analysis. Figure 1(a) shows a typical baseline subtracted waveform, recorded with PS, measured using an ${}^{241}\mathrm{Am}^{9}\mathrm{Be}$ neutron source. The rise time of the scintillation signal is around 2 ns and the waveform acquisition window is kept at 300 ns to capture the entire charge deposition from the ionizing particle. For the determination of the PSD, an integrated charge from short gate (SG) and the long gate (LG) are used which is described in detail later in the text. Normalized neutron and $\gamma$ waveforms, shown in Fig. 1(b), where the differences in the charge distribution between the two signals is observed at the tail region beyond 100ns. The inset in Fig. 1 (b) shows the zoomed version of the waveform from neutron and $\gamma$ event in the time window of 80 to 130 ns, to emphasis the differences in the waveforms for the PSD variable calculation.
@@
The radioactive $\gamma$ sources are used to study the response of the PS. The integrated charge (ADC) spectra from the ${}^{137}\mathrm{Cs}$ and ${}^{22}\mathrm{Na}$ radioactive sources are shown in the Fig. 2 (a) and (b), respectively. The prominent Compton edges can be seen indicating the energy deposition of the $\gamma$-rays in the PS. The PS detector is calibrated using the $\gamma$s from ${}^{137}\mathrm{Cs}$ and ${}^{22}\mathrm{Na}$ sources along with the 4.4 MeV $\gamma$ from an AmBe neutron source. The lowest measured ADC value used in our analysis is 200 which corresponds to a charge deposition of 151 $\mathrm{keV_{ee}}$. For the analysis, we have used the energy deposition in terms of ADC values only and not converted them in electron equivalent energy.
The neutron and $\gamma$ data are collected using an AmBe neutron source, kept at a distance of 5cm from the PS detector. No shielding or moderator is used between source and the PS. The neutrons emanating from the AmBe source range from 1 MeV to 12 MeV, peaked around 3-4 MeV. The neutron deposits energy in the PS by proton recoil and the energy deposition is quenched due to the electron density in the PS [20]. On the other hand, a 4.4 MeV $\gamma$ is emitted from the AmBe source produces the Compton scattered electron in the PS. The energy deposition, represented in the waveform, of the proton recoil and the Compton scattered electron is measured from the anode signal of the PMT and digitized using a waveform digitizer.
Figure 2: Integrated charge distribution of (a) ${}^{137}\textbf{Cs}$ and (b) ${}^{22}\textbf{Na}$ radioactive source.
@@ 3 Waveform Analysis
Organic scintillators exhibit fast and slow components due to their decay modes. The different ratios of fast and slow signals coming from various particles cause the pulses to have different shapes. Plastic scintillators have a very fast rise time of the $\sim$ 2 ns and decay time around $\sim$5 ns [21]. The interaction of $\gamma$ with organic scintillator is mediated by electromagnetic interaction, hence, having a fast rise and decay time. As for neutrons, due to collisional interaction mechanisms, neutrons have longer decay times. This difference in the neutron and $\gamma$ waveform from the EJ-276D PS is shown in Fig. 1 (b). The traditional CCM exploits the timing nature of different particles. Considering two distinct time windows SG and LG, charge is integrated to obtain $Q_{S}$ and $Q_{L}$, respectively. The ratio of $Q_{S}$ to $Q_{L}$, hence enables the pulse shape discrimination capability of the detector. The PSD variable, $PSD=1-\frac{Q_{S}}{Q_{L}}$, is calculated from each waveform and studied as a function of the total charge deposition in scintillator for the discrimination of neutron and $\gamma$ signals. Figure of Merit (F.o.M), described as $F.o.M=\frac{|\mu_{\gamma}-\mu_{n}|}{FWHM_{\gamma}+FWHM_{n}}$, is often used to estimate the degree of discrimination between the different particles. The $\mu_{\gamma}$ and $\mu_{n}$ represent the mean values of the neutron and $\gamma$ PSD distribution and $FWHM_{\gamma}$ and $FWHM_{n}$ represents the full width at half maximum of these distributions, respectively. To obtain a reasonable F.o.M, various combinations of $Q_{S}$ and $Q_{L}$ gate time windows are used in an iterative manner in the analysis. The optimized F.o.M of 1.21 is achieved for the PS using the integrated charge from the SG time window 60 to 108 ns and the LG time window 60 to 250 ns in the ADC range 1500 to 4000, where $\gamma$ and neutron events are clearly separable. For the same gate settings, the F.o.M in charge range 2000 to 4000 is found to be 1.35 implying higher ADC ranges have better neutron $\gamma$ discrimination. The SG and LG time windows mentioned above are fixed through out the analysis. Figure 3 (a) shows the PSD as a function of integrated charge measured from the AmBe neutron source. A clear separation between neutron and $\gamma$ events is observed above 1500 in the integrated charge range. The quantitative projection of PSD distributions are shown in Fig. 3 (b) and (c) for the lower charge and the higher charge ranges, respectively. It can be seen from Fig. 3 (b), that at lower charge values, between 200 to 800, the separation using PSD variable between neutron and $\gamma$ events worsens and the discrimination becomes difficult. Low signal to noise ratio, disparity in the Compton interaction, similar timing profiles of neutrons and $\gamma$s, resulting in effective failure of CCM in this region [22]. For the higher charge range (1500 to 4000), as shown in Fig. 3 (c), simple PSD based selection criteria can achieve a very good separation efficiency for discrimination of neutron and $\gamma$ events.
@@ 4.1 Multivariate Analysis
As an alternate to PSD technique, machine learning algorithms are being explored in various experiments [23, 24, 25] to distinguish different backgrounds. These methods do not directly rely on predefined charge integration windows, but can be trained to learn non-linear features from the raw waveforms. In this work, we will focus on the use of artificial neural networks (ANN), especially the MultiLayer Perceptron (MLP) [26] using a Bayesian inference, and a Support Vector Machine (SVM) [27] algorithms for the separation of neutron and $\gamma$ events.
@@
The MLP Bayesian Neural Network (MLPBNN) approach combines the structure of a traditional neural network with probabilistic modeling. In this framework, weights and biases are treated as probability distributions rather than fixed values. The network consists of one input layer, one or more hidden layers and a output layer. Bayesian inference is used to estimate these distributions, allowing the model to capture uncertainty in predictions. This approach improves generalization, offers uncertainty estimates, and reduces overfitting, especially in small or noisy datasets. MLPBNN is an efficient technique to tag the nonlinear correlation with associated uncertainties. In parallel, SVM algorithm which is also a supervised ML algorithm, used for classification and regression tasks. It works by finding the optimal hyperplane that best separates data points of different classes in a higher-dimensional space. The key idea is to maximize the margin between the closest data points (support vectors) and the hyperplane. SVMs can also use kernel functions to handle non-linear data by mapping it into a higher-dimensional space where a linear separator is possible. SVMs are known for their accuracy and effectiveness in high-dimensional spaces.
@@
While both methods have their advantages, MLPBNN needs fine-tuning for the involved hidden layers. Training MLPBNNs often takes more iterations to converge compared to traditional MLPs, particularly when using complex approximate inference techniques. Bayesian methods introduce additional hyperparameters (e.g., prior distributions, variational parameters) that require careful tuning, which can be non-trivial. On other hand, the SVM can be at many times slow to train and predict with large datasets, due to their complexity and memory requirements. The performance of SVMs heavily depends on the choice of kernel and parameters. They also do not provide direct estimation of probabilities for the predictions. For our studies we are using the TMVA toolkit [28], where the MLPBNN and SVM algorithms are readily available for the use. The advantage of using TMVA is the ease of integration of the packages to the ROOT based software from CERN [29], which enables the easy data handling, computation and plotting the results. To identify the optimal configuration, for both MLPBNN and SVM algorithms, optimization is done by tuning the set of input parameter dependencies and stability of the algorithms on the variables of interest. For e.g. for MLPBNN, we have tested the convergence of the algorithm with various hidden layers and epochs as well as optimizing the degree of correlation between the variables used for training. For SVM algorithm, the tolerance and $\Gamma$ values are studied systematically to obtain reasonable results.

### v2

@@ Abstract
In this work, we present results for discrimination of neutron and $\gamma$ events using a plastic scintillator detector with pulse shape discrimination capabilities. Machine learning (ML) algorithms are used to improve the discriminatory power between neutron and $\gamma$ events at lower energy ranges which otherwise are not addressed by the conventional pulse shape discrimination techniques. The use of a multilayer perceptron with Bayesian inference (MLPBNN) and support vector machine (SVM) algorithms are studied using the recorded waveforms from the detector. Input variables are constructed for the ML algorithms, which captures the essence of the differences in the head and tail part of the neutron and $\gamma$ waveforms. A new variable, which utilizes the product of kurtosis and variance calculated from the waveform gives better ranking in terms of separation of neutron and $\gamma$ events. The training and the testing of the ML algorithms are done using an AmBe neutron source. In the lower energy region, the results obtained from the ML predictions are compared with the results obtained from a time of flight (ToF) technique to benchmark the overall performance of the ML algorithms. A reasonable agreement is observed between the results obtained from ML algorithm and the ToF experiment in the studied energy range. The MLPBNN gives better discriminatory power for the neutron and $\gamma$ events than the SVM algorithm.
@@ 1 Introduction
Pulse Shape Discrimination (PSD) is a widely used technique for the particle identification in nuclear and particle physics experiments [nuclPSD, ProsJINST, junoPSD], as well as in applications such as reactor environment monitoring [prospNIMBack] and the national security [homeSec]. Among the various PSD methodologies, the charge comparison method (CCM) is one of the most reliable techniques for categorizing different types of ionizing radiation.
@@
Liquid organic scintillators are well-suited for PSD applications due to their well-defined scintillation time profiles for different particle interactions. The CCM effectively enhances these differences, making it a go-to method for particle identification in particle physics experiments. In particular, the use of this technique in reactor based neutrino detection experiments and rare event decay searches [DayaBayNIM, PROSPECT2019, ieee], where precise discrimination of neutron and $\gamma$ signals are required, to enhance the signal to background strength. In a very short baseline reactor based experiment, where both natural radioactivity and neutron-induced backgrounds (including fast and thermal neutrons) are prevalent, CCM provides a means of separating neutron-induced events from the actual signal of interest. Recently, developments in the plastic scintillator domain are being carried out to improvise the PSD properties [POZZI201319, ZAITSEVA201897]. Easy to handle, large scale production and modularity are some of the advantages of plastic scintillators and can be efficiently used as core detector for the anti-neutrino detection [SoLiD, PKN2022] and veto systems for muons and neutrons.
@@
Despite its advantages, the performance of CCM degrades significantly at lower energies, posing a major challenge for experiments demanding high background rejection sensitivities. As an example, the anti-neutrino experiments at very short baseline, few meters from the reactor core, suffers from the fast neutron background inside the reactor complexes [TEXONO:2018fse, DayaBay:2016ggj]. With optimal shielding configurations, these fast neutron background rates are reduced but can not be completely suppressed [MULMULE2018104]. The detection of anti-neutrino events, in these reactor based experiments, is generally through the process of inverse beta decay (IBD), which produces a positron and a neutron. The positron, assigned as the prompt signal, deposits its energy instantaneously followed by annihilation with surrounding electrons emitting two $\gamma$-rays. The neutron after thermalization in the detector medium is captured on the capture agent (like H, Li, Gd nuclei). The de-excitation of the capture agent emanates $\gamma$-rays, which forms the delayed signal. On the contrary, the fast neutrons interact with the detector via proton recoil reaction which mimics the prompt event signature and eventually undergoes thermalization and later gets captured by producing a delayed event signature. This can lead to a potential misidentification of a fast neutron event as an anti-neutrino event, even with the optimal shielding configurations [DayaBay2012, RONI2025]. Apart from the fast neutron background from the reactor, high energy neutrons, mostly from cosmic or muon induced production, may produce single or multiple neutrons while interacting with the detector material which usually contains carbon [ROGERS2019162436]. These secondary neutrons in turn may act as a potential source of additional neutron background in the experiments. Therefore, a comprehensive understanding of fast neutron background is vital for accurate event classification to improve the fidelity of neutrino experiments. The PSD technique is also extensively used in low energy nuclear physics experiments, especially in prompt fission neutron multiplicity measurements [PFNeutrons], can also benefit from these PSD based plastic scinitillators for neutron and $\gamma$ events.
[…]
@@ 2 Experimental Setup
A Plastic Scintillator (PS), EJ276D with PSD property from Eljen technology [eljen276d], was used to discriminate the signals from neutron and $\gamma$ events. The scintillator is of dimension $\phi$$\times51$ mm and has a maximum emission wavelength of 425 nm. The detector was optically coupled with a $\phi$51 mm diameter Hamamatsu R7724, 10 stage photo-multiplier tube (PMT) of borosilicate glass window, which was operated at the negative bias voltage of 1200 V. The PSD ability of this particular PS were studied before and a decent neutron and $\gamma$ discrimination was obtained for higher energies [Pant2024, Pagano2024]. The data acquisition system (DAQ), used to measure the anode signal from the PS, consists of a 8-channel CAEN V1751 digitizer [CAEN2024]. On-board digitization of the anode signal from the PMT was done using a leading edge discrimination technique. Triggered data was read by a Flash ADC of 10-bit resolution at 1 GSPS sampling rate. Raw waveforms were recorded and transferred to the computer via optical link for the detailed offline data analysis. Figure 1(a) shows a typical baseline subtracted waveform, recorded with PS, measured using an $^{241}\mathrm{Am}^{9}\mathrm{Be}$ neutron source. The rise time of the scintillation signal is around 2 ns and the waveform acquisition window is kept at 300 ns to capture the entire charge deposition from the ionizing particle. For the determination of the PSD, an integrated charge from short gate (SG) and the long gate (LG) are used which is described in detail later in the text. Normalized neutron and $\gamma$ waveforms, shown in Fig. 1(b), where the differences in the charge distribution between the two signals is observed at the tail region beyond 100 ns. The inset in Fig. 1 (b) shows the zoomed version of the waveform from neutron and $\gamma$ event in the time window of 80 to 130 ns, to emphasis the differences in the waveforms for the PSD variable calculation.
@@
[…]
The neutron and $\gamma$ data are collected using an AmBe neutron source, kept at a distance of 5cm from the PS detector. No shielding or moderator is used between source and the PS. The neutrons emanating from the AmBe source range from 1 MeV to 12 MeV, peaked around 3-4 MeV. The neutron deposits energy in the PS by proton recoil and the energy deposition is quenched due to the electron density in the PS [Reichhart:2011gz]. On the other hand, a 4.4 MeV $\gamma$ is emitted from the AmBe source produces the Compton scattered electron in the PS. The energy deposition, represented in the waveform, of the proton recoil and the Compton scattered electron is measured from the anode signal of the PMT and digitized using a waveform digitizer.
Figure 2: Integrated charge distribution of (a) $^{137}\textbf{Cs}$, (b) $^{22}\textbf{Na}$ radioactive source and (c) calibration obtained from the measured Compton edge energy deposition in the detector.
@@ 3 Waveform Analysis
Organic scintillators exhibit fast and slow components due to their decay modes. All charged particles interaction with the scintillator medium produces prompt fluorescence by exiting the singlet states. This process results in the decay time, in the order of a few nanoseconds, called a fast-time component. On the other hand, heavy charge particles, like proton and alpha particles, create higher ionization density of excited molecules in the scintillator medium as compared to electrons. This in turn, increases the probability of excitation of the triplet states leading to the delayed photon emission, called the slow component. The different ratios of fast and slow signals coming from various particles cause the pulses to have different shapes and enables the basis of PSD. Plastic scintillators have a very fast rise time of the $\sim$2 ns and decay time around $\sim$5 ns [knoll2010radiation].
@@
The interaction of $\gamma$ with organic scintillator is mediated by electromagnetic interaction, hence, having a fast rise and decay time. For the fast neutrons the dominant mechanism of energy transfer is through proton recoils, producing higher ionization density than electrons from $\gamma$-ray interactions. Due to the heavier mass of the proton, it loses energy slowly in the detector medium in comparison to electron which is generated by Compton interaction of $\gamma$-rays. Also the scintillation response time is slow for the protons due to the quenching mechanism [Awe_2021]. This results in a comparative longer tail part for the neutron waveforms. This difference in the neutron and $\gamma$ waveform from the EJ-276D PS is shown in Fig. 1 (b). The traditional CCM exploits the timing nature of different particles. Considering two distinct time windows SG and LG, charge is integrated to obtain $Q_{S}$ and $Q_{L}$, respectively. The ratio of $Q_{S}$ to $Q_{L}$, hence enables the pulse shape discrimination capability of the detector. The PSD variable, $PSD=1-\frac{Q_{S}}{Q_{L}}$, is calculated from each waveform and studied as a function of the total charge deposition in scintillator for the discrimination of neutron and $\gamma$ signals. Figure of Merit (F.o.M), described as $\mathrm{F.o.M=\frac{|\mu_{\gamma}-\mu_{n}|}{FWHM_{\gamma}+FWHM_{n}}}$, is often used to estimate the degree of discrimination between the different particles. The $\mathrm{\mu_{\gamma}}$ and $\mathrm{\mu_{n}}$ represent the mean values of the neutron and $\gamma$ PSD distribution and $\mathrm{FWHM_{\gamma}}$ and $\mathrm{FWHM_{n}}$ represents the full width at half maximum of these distributions, respectively. To obtain a reasonable F.o.M, various combinations of $Q_{S}$ and $Q_{L}$ gate time windows are used in an iterative manner in the analysis.
The optimized SG and LG corresponds to a time window of 60-108 ns and 60-250 ns, respectively. In the integrated charge range 1500-4000, a F.o.M of 1.21 is achieved, where the PSD discrimination still suffers from the non uniform energy loss by the proton recoils from neutron interactions in the detector. At low deposited energies, the number of detected photons is small, and statistical fluctuations in the fast and slow components reduce the neutron-$\gamma$ discrimination. In contrast, the F.o.M in the integrated charge range 2000-4000 improves to 1.35, implying when the energy of the particle increases, more photons are generated, which reduces the statistical fluctuations and both the fast component and slow component increase proportionally assuming the ionization density remains the same in both cases. Due to this, at higher charge region, the PSD is constant as a function of deposition energy in the measured range [sym15010011]. This shows the intrinsic pulse shape characteristic of a particle is constant, and the overall pulse height is proportional to the total deposited energy.
[…]
@@ 4.1 Multivariate Analysis
As an alternate to PSD technique, machine learning algorithms are being explored in various experiments [Yun:2024fxv, Yu:2025lxz, BOREXINO:2023nsa] to distinguish different backgrounds. These methods do not directly rely on predefined charge integration windows, but can be trained to learn non-linear features from the raw waveforms. In this work, we will focus on the use of artificial neural networks (ANN), especially the MultiLayer Perceptron (MLP) [Bourlard1994] using a Bayesian inference, and a Support Vector Machine (SVM) [svm] algorithms for the separation of neutron and $\gamma$ events.
@@
The MLP Bayesian Neural Network (MLPBNN) approach combines the structure of a traditional neural network with probabilistic modeling. In this framework, weights and biases are treated as probability distributions rather than fixed values. The network consists of one input layer, one or more hidden layers and a output layer. Bayesian inference is used to estimate these distributions, allowing the model to capture uncertainty in predictions. This approach improves generalization, offers uncertainty estimates, and reduces overfitting, especially in small or noisy datasets. MLPBNN is an efficient technique to tag the nonlinear correlation with associated uncertainties. This algorithm has been proven to be efficient for various practical application such as pattern recognition in image analysis [VEHTARI20001183]. In parallel, SVM algorithm which is also a supervised ML algorithm, used for classification and regression tasks. It works by finding the optimal hyperplane that best separates data points of different classes in a higher-dimensional space. The key idea is to maximize the margin between the closest data points (support vectors) and the hyperplane. SVMs can also use kernel functions to handle non-linear data by mapping it into a higher-dimensional space where a linear separator is possible. SVMs are known for their accuracy and effectiveness in high-dimensional spaces.
@@
While both methods have their advantages, MLPBNN needs fine-tuning for the involved hidden layers. Training MLPBNNs often takes more iterations to converge compared to traditional MLPs, particularly when using complex approximate inference techniques. Bayesian methods introduce additional hyperparameters (e.g., prior distributions, variational parameters) that require careful tuning, which can be non-trivial. On other hand, the SVM can be at many times slow to train and predict with large datasets, due to their complexity and memory requirements. The performance of SVMs heavily depends on the choice of kernel and parameters. They also do not provide direct estimation of probabilities for the predictions. For our studies we are using the TMVA toolkit [hoecker2009tmva], where the MLPBNN and SVM algorithms are readily available for the use. The advantage of using TMVA is the ease of integration of the packages to the ROOT based software from CERN [root], which enables the easy data handling, computation and plotting the results.